# Large-scale modality-invariant foundation models for brain MRI analysis: Application to lesion segmentation

### v1

@@ 2.5 Evaluation Setup
Downstream lesion segmentation was evaluated on three cohorts: (i) ATLAS v2 [10], a multi-center dataset of 655 post-stroke cases with $T_{1}$-weighted images and expert-annotated lesion masks; (ii) ISLES 2022 [11], a two-center dataset of 255 ischemic stroke cases with annotations on DWI and accompanying ADC and FLAIR sequences; and (iii) FCD BOHN [12], a single-center dataset of 85 focal cortical dysplasia (FCD) patients with manual lesion annotations in FLAIR and additional $T_{1}$-weighted MRI. All images were rigidly co-registered and pre-processed as FOMO60k.
Unique models were trained per imaging modality and tested on held-out sets using the Dice similarity coefficient (DSC). When feasible, test sets were constructed via center-wise splits; 10 randomly selected centers for ATLAS v2 (n = 255) and center #2 (n = 52) for ISLES 2022. FCD BOHN distinguishes a subset of 28 challenging cases, which served as the held-out set. Given the absence of additional modalities in ATLAS v2, the training set (n = 400) was further partitioned into subsets of 100%, 75%, 50%, and 25% to assess robustness to reduced sample size. Results were compared against training from scratch (i.e., no pre-training) baselines.
@@ 3.2 Lesion Segmentation Performance
Fig. 3: Segmentation DSC in the BOHN FCD and ISLES 2022 datasets for each modality and pre-training method.

### v2

@@ 2.5 Evaluation Setup
Downstream lesion segmentation was evaluated on three cohorts: (i) ATLAS v2 [7], a multi-center dataset of 655 post-stroke cases with $T_{1}$-weighted images and expert-annotated lesion masks; (ii) ISLES 2022 [5], a two-center dataset of 255 ischemic stroke cases with annotations on DWI and accompanying ADC and FLAIR sequences; and (iii) FCD BONN [9], a single-center dataset of 85 focal cortical dysplasia (FCD) patients with manual lesion annotations in FLAIR and additional $T_{1}$-weighted MRI. All images were rigidly co-registered and pre-processed as FOMO60k.
Unique models were trained per imaging modality and tested on held-out sets using the Dice similarity coefficient (DSC). When feasible, test sets were constructed via center-wise splits; 10 randomly selected centers for ATLAS v2 (n = 255) and center #2 (n = 52) for ISLES 2022. FCD BONN distinguishes a subset of 28 challenging cases, which served as the held-out set. Given the absence of additional modalities in ATLAS v2, the training set (n = 400) was further partitioned into subsets of 100%, 75%, 50%, and 25% to assess robustness to reduced sample size. Results were compared against training from scratch (i.e., no pre-training) baselines.
@@ 3.2 Lesion Segmentation Performance
Fig. 3: Segmentation DSC in the BONN FCD and ISLES 2022 datasets for each modality and pre-training method.